OpenStudy (scarlettfarra2000):

Would someone be so kind to check to see if this is right please? Simplify the expression (Simplify your answer. Use positive exponents only.)

OpenStudy (scarlettfarra2000):

\[\frac{ 7q ^{-1}k ^{-3} }{ w ^{0} }\] Is the problem

OpenStudy (scarlettfarra2000):

For every nozero number a, \[a ^{0}=1\] using this property to rewrite the denominator \[\frac{ 7q ^{-1}k ^{-3} }{ 1}\]

OpenStudy (mathstudent55):

So far, you're correct. You can simplify more since \(\dfrac{a}{1} = a\) for every \(a\).

OpenStudy (scarlettfarra2000):

For every nozero number a and b integer n, \[a ^{-n}=\frac{ 1 }{ a ^{n}}\] \[7q ^{-1}k ^{-3}=7\left(\begin{matrix}1 \\ q\end{matrix}\right)k ^{-3}\]

OpenStudy (mathstudent55):

Definition of a negative exponent: For every nonzero \(a\), \(a^{-n} = \dfrac{1}{a^n} \)

OpenStudy (scarlettfarra2000):

Using this property \[p ^{-1}\] and \[k ^{-3}\]\[7(\frac{ 1 }{ q }) (\frac{ 1 }{ k ^{3} })\]

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now you have positive exponents. You can combine the entire thing into one single fraction.

OpenStudy (scarlettfarra2000):

\[\frac{ 7 }{ qk ^{3} }?\]

OpenStudy (mathstudent55):

Correct.

OpenStudy (scarlettfarra2000):

Thank you @mathstudent55

OpenStudy (mathstudent55):

Summarizing the entire process: \(\dfrac{ 7q ^{-1}k ^{-3} }{ w ^{0} } =\) \(= \dfrac{ 7q ^{-1}k ^{-3} }{ 1 }\) \(= \dfrac{ 7}{q k ^{3} }\)

OpenStudy (mathstudent55):

You're welcome. In general, a positive exponent in the numerator is a negative exponent in the denominator, and vice-versa.

OpenStudy (mathstudent55):

Example: